OpenStudy (anonymous):

A spherical object has a diameter of 18 cm. It has a spherical inner core with a diameter of 9 cm. What is the volume of the other layer? Use 3.14 to approximate pi and express your answer in hundredths.

OpenStudy (academicgurusinc):

Volume of a sphere = 4/3 pi r^2 First find the volume of the outer sphere. V = 4/3 * pi * (0.09 ^2) = 0.034 m^3 Next find the volume of the inner speher V = 4/3 * pi * (0.045^2) = 0.0085 To find the volume of the other layer, subtract the volume of the core from the volume of the Outer surface. V = 0.034 - 0.0085 = 0.0255 ~ 0.03 m^3

OpenStudy (academicgurusinc):

I hope that helps. If you found this helpful, I encourage you to subscribe to our youtube channel, to stay current with all of our new videos: https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg You may also want to check out our Math Challenge contest, for a chance to win a $50 Visa giftcard: https://www.youtube.com/watch?v=2Dbu-R_Hj7E Regards, Academic Gurus Inc.

OpenStudy (anonymous):

0.03